# Chirally motivated $\pi\Sigma-\bar{K}N$ model in a finite volume

### v1

@@ 2 Finite volume model formulation
where $l^{2}\equiv|\vec{l}|^{2}$. We need to generalize Eq. (2) to the case of a cubic FV $V=L^{3}$. For a free particle inside a finite cube of side length $L$, with periodic boundary conditions, the possible three-momenta are $\vec{p}_{n}=\frac{2\pi}{L}\vec{n}$, $\vec{n}\in\mathds{Z}^{3}$. The discretization of the integral in Eq. (2) is straightforward and yields
@@
where $n=|\vec{n}|$, the $\pm$ sign in the exponent on the last line matches the sign of the imaginary part of the momentum $q$, and the tilded $G$-functions stand for the loop function parts divided by the form-factor $g(q)$ squared, e.g. $G_{P}(q)=[g(q)]^{2}\tilde{G}_{\rm P}(q)$. The terms $\Delta\tilde{G}_{\cal R}(q;L)$ and $\Delta\tilde{G}_{\cal I}(q;L)$ represent FV corrections to the first (pseudo-real) and second (phase-space) parts of the infinite volume loop function, Eq. (2), and apparently vanish in the $L\rightarrow\infty$ limit. We also note that, for a fixed $L$-value, the oscillations caused by the exp(±$(\pm{\rm i}qLn)$ term in $\Delta\tilde{G}_{\cal I}(q;L)$ are also better suppressed at complex energies further from the real axis.
@@ 3 Results and discussion
In principle, we could also tune the Prague model parameters to fit the BaSc energy spectra generated for the irreps shown in BaSc:2023ori but it would be a more involved task. In the present work we just demonstrate how our finite-volume model does in comparison with the BaSc energy spectrum of the $G_{1u}(0)$ irrep that seems most relevant when one considers only the s-wave. In Table 1 we present the stationary state energies generated by our FV model for two parameter settings, P0 and P2 taken from Bruns:2021krp ; Cieply:2023saa , and the BaSc energy spectrum of the $G_{1u}(0)$ irrep extracted from Fig. 8 in BaSc:2023ori . The P0 spectrum matches exactly the energies of divergencies observed in the bottom panel of Fig. 2 and the energy variations shown for the P0 and P2 spectra were obtained by varying the spatial extent by $\Delta L=1$ GeV-1, a value about 3 times larger than the standard deviation of the $L$-value adopted in BaSc:2023ori . The two lowest energies (both below the $\bar{K}N$ threshold) appear particularly stable with respect to the $L$-variations in our model. It should be noted that the P2 model does not reproduce the $K^{-}p$ reactions data so well as the P0 model does due to complimenting the model fit with the $\pi\Sigma$ photoproduction data. The P2 parameter setting is also closely related to a specific construction of the photoproduction kernel, see Cieply:2023saa for details, and does not provide realistic predictions at higher energies, e.g. for the $\Lambda(1670)$ pole position. Therefore, the P0 model results should be regarded more seriously while the P2 model spectrum of FV stationary energies was included in Table 1 just to demonstrate a possible model dependence. As one can see, the correspondence between the P0 and BaSc energy spectra is remarkably good with only one energy being off by more than one standard deviation reported for the $G_{1u}(0)$ irrep. We would like to emphasize the point that these two energy spectra result from two very different concepts, one being the $\pi\Sigma-\bar{K}N$ coupled channel model based on chiral dynamics with parameters fitted to real physics data and the other the LQCD simulations of the MB interactions. It is really encouraging to see that so nice mutual agreement can be reached between these two distinctly different worlds.
@@ Appendix A Loop function in a finite volume
$G^{\rm FV}_{\cal R}(q;L)$ $=4\pi[g(q)]^{2}(\alpha^{2}-q^{2})\int\frac{d^{3}l}{(2\pi)^{3}}\left(\frac{2\pi}{L}\right)^{3}\sum_{\vec{n}\in\mathds{Z}^{3}}\frac{\delta^{3}(\vec{l}-\vec{p}_{n})}{(l^{2}+\alpha^{2})^{2}}$
=$\overset{\mathrm{PSF}}{=}4\pi[g(q)]^{2}(\alpha^{2}-q^{2})\int\frac{d^{3}l}{(2\pi)^{3}}\sum_{\vec{k}\in\mathds{Z}^{3}}\frac{e^{{\rm i}L(\vec{k}\cdot\vec{l})}}{(l^{2}+\alpha^{2})^{2}}$
$=[g(q)]^{2}\frac{\alpha^{2}-q^{2}}{2\alpha}\sum_{\vec{k}\in\mathds{Z}^{3}}e^{-\alpha L|\vec{k}|}$
$=[g(q)]^{2}\frac{\alpha^{2}-q^{2}}{2\alpha}\left(1+6e^{-\alpha L}+12e^{-\sqrt{2}\alpha L}+8e^{-\sqrt{3}\alpha L}+\ldots\right)\;.$ (8)
@@
$G^{\rm FV}_{\cal I}(q;L)$ $=4\pi[g(q)]^{2}\int\frac{d^{3}l}{(2\pi)^{3}}\left(\frac{2\pi}{L}\right)^{3}\sum_{\vec{n}\in\mathds{Z}^{3}}\delta^{3}(\vec{l}-\vec{p}_{n})\left[\frac{1}{l^{2}-q^{2}}-\frac{l^{2}+3\alpha^{2}}{(l^{2}+\alpha^{2})^{2}}\right]$
@@
=$[g(q)]^{2}\left[\pm{\rm i}\,q+\sum_{\vec{0}\not=\vec{k}\in\mathds{Z}^{3}}\frac{1}{|\vec{k}|L}\left(e^{\pm{\rm i}qL|\vec{k}|}-(1+\alpha L|\vec{k}|)e^{-\alpha L|\vec{k}|}\right)\right]\;,$ (9)

### v2

@@ 2 Finite volume model formulation
where $l^{2}\equiv|\vec{l}|^{2}$. Despite its simple non-relativistic appearance, our loop function has the correct relativistic absorptive part $\sim{\rm i}q$ due to the use of relativistic kinematics, see Eq. (10) in Appendix B. The form factors $g(l)$ regulate the propagation of particles with momenta $l\lesssim\alpha$ and have provided a convenient tool for extrapolations of the MB amplitudes off-the-energy shell since their introduction when the adopted coupled-channel model was designed in Kaiser:1995eg . A comparison with an alternative of dimensionally regularized loop function was given e.g. in Bruns:2022sio , see the discussion following Eq. (2.11) there.
We need to generalize Eq. (2) to the case of a cubic FV $V=L^{3}$. For a free particle inside a finite cube of side length $L$, with periodic boundary conditions, the possible three-momenta are $\vec{p}_{n}=\frac{2\pi}{L}\vec{n}$, $\vec{n}\in\mathds{Z}^{3}$. The discretization of the integral in Eq. (2) is straightforward and yields
@@
where $n=|\vec{n}|$, the $\pm$ sign in the exponent on the last line matches the sign of the imaginary part of the momentum $q$, and the tilded $G$-functions stand for the loop function parts divided by the form-factor $g(q)$ squared, e.g. $G_{P}(q)=[g(q)]^{2}\tilde{G}_{\rm P}(q)$. The terms $\Delta\tilde{G}_{\cal R}(q;L)$ and $\Delta\tilde{G}_{\cal I}(q;L)$ represent FV corrections to the first (pseudo-real) and second (phase-space) parts of the infinite volume loop function, Eq. (2), and apparently vanish in the $L\rightarrow\infty$ limit. We also note that, for a fixed $L$-value, the oscillations caused by the exp(±$(\pm{\rm i}qLn)$ term in $\Delta\tilde{G}_{\cal I}(q;L)$ are better suppressed at complex energies further from the real axis.
@@ 3 Results and discussion
In principle, we could also tune the Prague model parameters to fit the BaSc energy spectra generated for the irreps shown in BaSc:2023ori but it would be a more involved task. In the present work we just demonstrate how our finite-volume model does in comparison with the BaSc energy spectrum of the $G_{1u}(0)$ irrep that seems most relevant when one considers only the s-wave. In Table 1 we present the stationary state energies generated by our FV model for two parameter settings, P0 and P2 taken from Bruns:2021krp ; Cieply:2023saa , and the BaSc energy spectrum of the $G_{1u}(0)$ irrep extracted from Fig. 8 in BaSc:2023ori . The P0 spectrum matches exactly the energies of divergencies observed in the bottom panel of Fig. 2 and the energy variations shown for the P0 and P2 spectra were obtained by varying the spatial extent by $\Delta L=1$ GeV-1, a value about 3 times larger than the standard deviation of the $L$-value adopted in BaSc:2023ori . The two lowest energies (both below the $\bar{K}N$ threshold) appear particularly stable with respect to the $L$-variations in our model. It should be noted that the P2 model does not reproduce the $K^{-}p$ reactions data so well as the P0 model does due to complementing the model fit with the $\pi\Sigma$ photoproduction data. The P2 parameter setting is also closely related to a specific construction of the photoproduction kernel, see Cieply:2023saa for details, and does not provide realistic predictions at higher energies, e.g. for the $\Lambda(1670)$ pole position. Therefore, the P0 model results should be regarded more seriously while the P2 model spectrum of FV stationary energies was included in Table 1 just to demonstrate a possible model dependence. As one can see, the correspondence between the P0 and BaSc energy spectra is remarkably good with only one energy being off by more than one standard deviation reported for the $G_{1u}(0)$ irrep. We would like to emphasize the point that these two energy spectra result from two very different approaches to QCD, one being the $\pi\Sigma-\bar{K}N$ coupled channel model based on chiral dynamics with parameters fitted to real physics data and the other the LQCD simulations of the MB interactions. It is really encouraging to see that so nice mutual agreement can be reached between these two distinctly different worlds.
@@ Appendix A Loop function in a finite volume
$\tilde{G}^{\rm FV}_{\cal R}(q;L)$ $=G^{\rm FV}_{\cal R}(q;L)/[g(q)]^{2}=4\pi(\alpha^{2}-q^{2})\int\frac{d^{3}l}{(2\pi)^{3}}\left(\frac{2\pi}{L}\right)^{3}\sum_{\vec{n}\in\mathds{Z}^{3}}\frac{\delta^{3}(\vec{l}-\vec{p}_{n})}{(l^{2}+\alpha^{2})^{2}}$
=$\overset{\mathrm{PSF}}{=}4\pi(\alpha^{2}-q^{2})\int\frac{d^{3}l}{(2\pi)^{3}}\sum_{\vec{k}\in\mathds{Z}^{3}}\frac{e^{{\rm i}L(\vec{k}\cdot\vec{l})}}{(l^{2}+\alpha^{2})^{2}}=\frac{\alpha^{2}-q^{2}}{2\alpha}\sum_{\vec{k}\in\mathds{Z}^{3}}e^{-\alpha L|\vec{k}|}$
$=\frac{\alpha^{2}-q^{2}}{2\alpha}\left(1+6e^{-\alpha L}+12e^{-\sqrt{2}\alpha L}+8e^{-\sqrt{3}\alpha L}+\ldots\right)\;.$ (8)
@@
$\tilde{G}^{\rm FV}_{\cal I}(q;L)$ $=G^{\rm FV}_{\cal I}(q;L)/[g(q)]^{2}=4\pi\int\frac{d^{3}l}{(2\pi)^{3}}\left(\frac{2\pi}{L}\right)^{3}\sum_{\vec{n}\in\mathds{Z}^{3}}\delta^{3}(\vec{l}-\vec{p}_{n})\left[\frac{1}{l^{2}-q^{2}}-\frac{l^{2}+3\alpha^{2}}{(l^{2}+\alpha^{2})^{2}}\right]$
@@
=±$\pm{\rm i}\,q+\!\sum_{\vec{0}\not=\vec{k}\in\mathds{Z}^{3}}\frac{1}{|\vec{k}|L}\left(e^{\pm{\rm i}qL|\vec{k}|}-(1+\alpha L|\vec{k}|)e^{-\alpha L|\vec{k}|}\right)\;,$ (9)